# Quantifying Statistical Significance in Diffusion-Based Anomaly Localization via Selective Inference

### v4

@@ 4 Selective inference for diffusion-based anomaly localization
In this section, we introduce selective inference (SI) framework for testing the anomalous regions detected by a diffusion model and propose a method to perform valid hypotheses tests.
@@ 4.1 Computing valid $p$-values via selective inference
As mentioned in §3, we need to identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$ to compute the $p$-values. If we ignore the fact that the anomalous region is identified by a diffusion model, the test statistic in (11) is simply a linear transformation of the Gaussian random vectors $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$, and hence itself follows a Gaussian distribution:
$\begin{gathered}T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}),\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.\end{gathered}$ (14)
However, as mentioned in §3, in reality the dependence of $\bm{\nu}$ on $\bm{x}$ via the diffusion model is more intricate, making the sampling distribution intractably complex. Consequently, obtaining this sampling distribution directly is challenging.
@@
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The sufficient statistics of the nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
@@
The proof of Theorem 1 is given in Appendix A.1. The following theorem tells that the selective $p$-value can be analytically derived from the conditional sampling distribution, which follows a truncated Gaussian distribution.
@@
The proof of the Theorem 2 is given in Appendix A.2. Once the truncation intervals $\mathcal{Z}$ are identified, computing the selective $p$-value in (22) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
4.2 Identification of subinterval $\mathcal{Z}^{\mathrm{sub}}$ in $\mathcal{Z}$
To tackle the identification of the truncation invervals $\mathcal{Z}$, we employ a divide-and-conquer strategy. Directly characterizing $\mathcal{Z}$ is challenging due to the complexity of the diffusion model’s computation. Our method decomposes the $n$-dimensional data space into a collection of polyhedra by imposing additional conditioning, a process we refer to as over-conditioning (OC) Duy and Takeuchi (2022). Each polyhedron in the $n$-dimensional space corresponds to an interval in the one-dimensional space $\mathcal{Z}$. Thus, we can examine these intervals sequentially in $\mathcal{Z}$ to determine whether they yield the same selected anomalous regions as observed. To this end, we need to identify a subinterval $\mathcal{Z}^{\mathrm{sub}}\subseteq\mathcal{Z}$. We show that the anomalous region from a diffusion model can be characterized by a piecewise-linear mapping, and that for each $z\in\mathbb{R}$, the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed analytically by solving a system of linear inequalities (see Appendix B for details).
4.3 Identification of $\mathcal{Z}$
@@
An algorithm for computing the selective $p$-value is summarized in Algorithm 2. Figure 2 illustrates the example of conditional sampling distribution. It shows how the subintervals determine the conditional sampling distribution.
@@
Algorithm 2 Selective $p$-value by Parametric Programming
0:  $\bm{x},\bm{x}^{\mathrm{ref}}$
1:  Initialize $\mathcal{Z}\leftarrow\emptyset$
2:  Compute $\mathcal{M}_{\bm{x}}$, $\bm{a}$, $\bm{b}$, and $T(\bm{x},\bm{x}^{\mathrm{ref}})$ by (8), (21), (11)
3:  Initialize $z$ to a sufficiently small value
4:  while $z$ is not large enough do
5:     Compute $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (34)
6:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}$ then
7:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$
8:     end if
9:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)}+\gamma$, where $0<\gamma\ll 1$
10:  end while
11:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}\right)$
11:  $p_{\mathrm{selective}}$
@@ 5 Experiments
  • w/o-pp: An ablation study without the parametric programming technique described in §4.3. The $p$-value is computed using (22), with $\mathcal{Z}$ replaced by $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})$.
@@
The architecture of diffusion models used in all experiments is detailed in Appendix D. We executed all experiments on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
@@ 5.1 Numerical experiments
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. In the type I error rate experiments, we used only normal images. The synthetic dataset for normal images is generated to follow $\bm{X}\sim\mathcal{N}(\bm{0},\Sigma)$. We made 1,000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only abnormal images. We generated 1,000 abnormal images $\bm{X}\sim\mathcal{N}(\bm{\mu},\Sigma)$. The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen. The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$. In all experiments, we made the synthetic dataset for 1,000 reference images to follow $\bm{X}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma)$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$. All experiments were conducted under the significance level $\alpha=0.05$. The diffusion models were trained on the normal images from the synthetic dataset. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings (see Appendix E for details). The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In all experiment, we generated normal-looking images through probabilistic sampling, $\eta$ was set to $1$. In addition, we conducted robustness experiments against non-Gaussian noise. The details of the robustness experiments are described in Appendix F.
@@ 5.2 Real-world data experiments
We conducted experiments using T2-FLAIR MRI brain scans from the Brain Tumor Segmentation (BraTS) 2023 dataset Karargyris et al. (2023); LaBella et al. (2023) and MVtec AD dataset Bergmann et al. (2019). The details of the experimental settings are described in Appendix G.
@@ Results
Table 1: Type I error rate and power comparison on real-world datasets at the significance level of $\alpha(=0.05)$. The proposed and bonferroni methods control the type I error rate at or below $\alpha$ for each dataset, whereas naive fails to do so and is therefore unreliable. Each entry shows type I error rate (left) and power (right). Additionally, proposed achieves higher power than bonferroni. The Figures 9 and 10 show the results of the proposed and naive methods for the MRI images (BraTS) and MVTec AD dataset, respectively.
Dataset naive bonferroni proposed
Bottle 0.46/N.A 0.00/0.00 $\bm{0.04}$/$\bm{0.18}$
Cable 0.88/N.A 0.00/0.00 $\bm{0.02}$/$\bm{0.40}$
Grid 0.82/N.A 0.00/0.00 $\bm{0.06}$/$\bm{0.34}$
Transistor 0.86/N.A 0.00/0.00 $\bm{0.08}$/$\bm{0.28}$
BraTS (T2-FLAIR) 0.59/N.A 0.00/0.00 $\bm{0.05}$/$\bm{0.28}$
Table 1 shows the comparison of the type I error rate and power. The naive cannot control the type I error rate, while the proposed and bonferroni can control the type I error rate below the significance level $\alpha$. The proposed has higher power than bonferroni. In addition, the examples of the results for each image are shown in Appendix H.
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.527$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
(a) bottle (Left: Normal, Right: Anomaly)
@@
(b) Brain (Left: Normal, Right: Anomaly)
Figure 5: Real-world examples of the naive and proposed methods. For each sample, the $p_{\mathrm{selective}}$ is high (true negative) for normal images and low (true positive) for abnormal images. In contrast, the $p_{\mathrm{naive}}$ remains low for all images, indicating that it fails to control the type I error rate.
6 Conclusions, limitations and future works
We introduced the DAL-Test, a novel statistical procedure for anomaly localizations identified by a diffusion model. With the proposed DAL-Test, the false positive detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by using a diffusion model. We demonstrated that the DAL-Test has higher power than the bonferroni correction, the only existing method for controlling the false positive detection rate. However, the growing the size of the diffusion model also leads to increased computational demands. In future work, we will focus on improving the computational efficiency of the DAL-Test.
@@ A.1 Proof of Theorem 1
$p_{\mathrm{selective}}\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\sim\mathrm{Uniform}(0,1),$ (24)
@@
$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\right)=\alpha,\;\forall\alpha\in(0,1).$ (25)
@@
$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\right)$ (26)
$=\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}})\ \mathbb{P}_{\mathrm{H}_{0}}(\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}})d\mathcal{Q}_{\bm{x}}$
[…]
@@ A.2 Proof of Theorem 2
=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z\right\}$
=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\}$
=$=\left\{\bm{a}+\bm{b}z\;\middle|\;z\in\mathcal{Z}\right\},$
where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$. As a result

### v5

@@ 4 Selective inference for diffusion-based anomaly localization
In this section, we introduce the SI framework for testing the anomalous regions detected by a diffusion model and propose a method to perform valid hypothesis tests.
@@ 4.1 Computing valid $p$-values via selective inference
As mentioned in §3, we need to identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$ to compute the $p$-values. If we ignore the fact that the anomalous region is identified by a diffusion model, the test statistic in (11) is simply a linear transformation of the Gaussian random vectors $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$, and hence itself follows a Gaussian distribution under the null hypothesis $\mathrm{H}_{0}$:
[…]
where $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$ are assumed to be independent. However, as mentioned in §3, in reality the dependence of $\bm{\nu}$ on $\bm{x}$ via the diffusion model is more intricate, making the sampling distribution intractably complex. Consequently, obtaining this sampling distribution directly is challenging.
@@
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The sufficient statistic of the nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
@@
The proof of Theorem 1 is given in Appendix A.1. The following theorem states that the selective $p$-value can be analytically derived from the conditional sampling distribution, which follows a truncated Gaussian distribution.
@@
The proof of Theorem 2 is given in Appendix A.2. Once the truncation intervals $\mathcal{Z}$ are identified, computing the selective $p$-value in (22) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
4.2 Identification of truncation intervals $\mathcal{Z}$
Identification of subinterval $\mathcal{Z}^{\mathrm{sub}}$
To tackle the identification of the truncation intervals $\mathcal{Z}$, we employ a divide-and-conquer strategy. Directly characterizing $\mathcal{Z}$ is challenging due to the complexity of the diffusion model’s computation. Our method decomposes the $n$-dimensional data space into a collection of polyhedra by imposing additional conditioning, a process we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022). Each polyhedron in the $n$-dimensional space corresponds to an interval in the one-dimensional space $\mathcal{Z}$. Thus, we can examine these intervals sequentially in $\mathcal{Z}$ to determine whether they yield the same selected anomalous regions as observed. To this end, we need to identify a subinterval $\mathcal{Z}^{\mathrm{sub}}\subseteq\mathcal{Z}$. We show that the anomalous region from a diffusion model can be characterized by a piecewise-linear mapping. Exploiting this property, the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed analytically for each $z\in\mathbb{R}$ by solving a system of linear inequalities (see Appendix B).
Identification of $\mathcal{Z}$ via parametric programming
@@ 4.2 Identification of truncation intervals $\mathcal{Z}$
Figure 2 illustrates the conditional sampling distribution determined by these truncation intervals. An algorithm for computing the selective $p$-value is summarized in Algorithm 2.
Algorithm 2 Selective $p$-value by Parametric Programming
[…]
2:Initialize $\mathcal{Z}\leftarrow\emptyset$
3:Compute $\mathcal{M}_{\bm{x}}$, $\bm{a}$, $\bm{b}$, and $T(\bm{x},\bm{x}^{\mathrm{ref}})$ by (8), (11), (21)
4:Initialize $z$ to a sufficiently small value
5:while $z$ is not large enough do
6:  Compute $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (34)
7:  if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}$ then
8:   $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$
9:  end if
10:  $z\leftarrow\max{\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)}+\gamma$, where $0<\gamma\ll 1$
11:end while
[…]
@@ 5 Experiments
  • w/o-$p$: An ablation study without the parametric programming technique described in §4.2. The $p$-value is computed using (22), with $\mathcal{Z}$ replaced by $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})$.
@@
The architecture of diffusion models used in all experiments is detailed in Appendix D. We executed all experiments on AMD EPYC 9474F processor, 48-core 3.63.6GHz CPU and 768768GB memory.
@@ 5.1 Numerical experiments
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. In the type I error rate experiments, we used only normal images. The synthetic dataset for normal images is generated to follow $\bm{X}\sim\mathcal{N}(\bm{0},\Sigma)$. We generated 1,000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only abnormal images. We generated 1,000 abnormal images $\bm{X}\sim\mathcal{N}(\bm{\mu},\Sigma)$. The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen. The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$. In all experiments, we generated the synthetic dataset for 1,000 reference images to follow $\bm{X}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma)$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$. All experiments were conducted under the significance level $\alpha=0.05$. The diffusion models were trained on the normal images from the synthetic dataset. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was performed using $5$-step sampling (see Appendix E for details). The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In all experiments, we generated normal-looking images through probabilistic sampling, with $\eta$ set to $1$.
@@ 5.2 Real-world data experiments
We conducted experiments using T2-FLAIR MRI brain scans from the Brain Tumor Segmentation (BraTS) 2023 dataset (Karargyris et al., 2023; LaBella et al., 2023) and MVTec AD dataset (Bergmann et al., 2019). The details of the experimental settings are described in Appendix F.
@@ Results
Table 1: Type I error rate and power comparison on real-world datasets at the significance level of $\alpha(=0.05)$. The proposed and bonferroni methods control the type I error rate at or below $\alpha$ for each dataset, whereas naive fails to do so and is therefore unreliable. Additionally, proposed achieves higher power than bonferroni. Figure 5 shows qualitative results of the proposed and naive methods on the MVTec AD dataset and the BraTS dataset.
naive bonferroni proposed
Dataset Type I Error Power Type I Error Power Type I Error Power
Bottle 0.46 N.A 0.00 0.00 $\bm{0.04}$ $\bm{0.18}$
Cable 0.88 N.A 0.00 0.00 $\bm{0.02}$ $\bm{0.40}$
Grid 0.82 N.A 0.00 0.00 $\bm{0.06}$ $\bm{0.34}$
Transistor 0.86 N.A 0.00 0.00 $\bm{0.08}$ $\bm{0.28}$
BraTS (T2-FLAIR) 0.59 N.A 0.00 0.00 $\bm{0.05}$ $\bm{0.28}$
Table 1 shows the comparison of the type I error rate and power. The naive cannot control the type I error rate, while the proposed and bonferroni can control the type I error rate below the significance level $\alpha$. The proposed has higher power than bonferroni.
@@
(a) Brain MRI (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.907$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.010$
(b) Bottle (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.567$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
(c) Cable (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.028$, $p_{\rm selective}:0.789$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
(d) Grid (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.769$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
(e) Transistor (Left: Normal, Right: Anomaly)
Figure 5: Results of the proposed and naive methods on the BraTS dataset (one example) and the MVTec AD dataset (four categories). For each dataset, the left image is a normal sample and the right image is an anomalous sample.
6 Conclusions
We introduced the DAL-Test, a novel statistical procedure for anomaly localization identified by a diffusion model. With the proposed DAL-Test, the false positive detection rate can be controlled at the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by using a diffusion model. We demonstrated that the DAL-Test has higher power than the Bonferroni correction, the only existing method for controlling the false positive detection rate.
@@ A.1 Proof of Theorem 1
$p_{\mathrm{selective}}\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\sim\mathrm{Uniform}(0,1),$ (24)
@@
$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\right)=\alpha,\;\forall\alpha\in(0,1).$ (25)
@@
$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\,\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right)$ (26)
$=\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}\!\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\,\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right)$
$\qquad\qquad\cdot\,\mathbb{P}_{\mathrm{H}_{0}}\!\left(\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}}\right)d\mathcal{Q}_{\bm{x}}$
$=\alpha\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}\!\left(\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}}\right)d\mathcal{Q}_{\bm{x}}$
@@ A.2 Proof of Theorem 2
$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z\right\}$
$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\}$
$=\left\{\bm{a}+\bm{b}z\;\middle|\;z\in\mathcal{Z}\right\},$
where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$. As a result,